A horizontal force of 250 N is applied to a stationary wooden Box in one direction and a 600 N horizontal force is applied in the opposite direction what additional force is needed for the Box to remain stationary
Answer by ikleyn(52798)

.

    250 + x = 600


          x = 600 - 250 = 350.


ANSWER.  The force of 350 N must be applied in parallel with the force of 250 N in the same direction.

Solved, answered and explained.